OpenStudy (anonymous):

1/4 (17=5n) Please show all work and simplify all the way :)

OpenStudy (anonymous):

really need to edit this one is it \[\frac{1}{4}(17+5n)\]?

OpenStudy (anonymous):

@satellite73 Yeah

OpenStudy (anonymous):

I made a few mistakes, sorry :(

OpenStudy (anonymous):

then there is no "simplifying" you can however multiply using the distributive law and get \[\large \frac{17}{4}+\frac{5}{4}n\]

OpenStudy (anonymous):

Thank You!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ask your math teacher (gently) why one form is "simpler" than another

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

...gently...

OpenStudy (anonymous):

ok don't